# Rapid Word Learning Through Meta In-Context Learning

### v2

@@ Abstract
Wentao Wang1     Guangyuan Jiang2     Tal Linzen1     Brenden M. Lake1 1New York University     2Peking University {ww2135, linzen, brenden}@nyu.edu     jgy@stu.pku.edu.cn
@@ 1 Introduction
We also finetune Llama-3 8B with Minnow to see if we can enhance its word-learning ability. In a series of discriminative and generative evaluations, we show that this improves Llama-3 8B’s ability to discriminate between new words, identify syntactic categories of new words, and generate reasonable new usages and definitions for new words, where each new word is learned from one or a few in-context examples. Most of these improvements are achieved without specific training on these evaluation tasks. We will release our code upon publication of our work.
@@ 3.1 Method: Minnow
Note that our setting differs from previous MetaICL settings (Min et al., 2022; Chen et al., 2022; Lake and Baroni, 2023) in two ways. First, each example is not an input–output pair $(x^{(i)}_{k},y^{(i)}_{k})$, but just $x^{(i)}_{k}$. Second, there is no explicit separation between study examples and a query: our setting effectively uses every example $x^{(i)}_{k}$ as a query with all previous examples $x^{(i)}_{1},\ldots,x^{(i)}_{k-1}$ as its study examples.
@@ 3.2 Data
We split both the meta-learning component (by word) and the language modeling component (by sentence) into training (80%), validation (10%) and test (10%) portions. Each dataset is used for both training models from scratch and finetuning pre-trained Llama-3 8B, but the text is formatted and tokenized differently (in addition to the different special tokens in Section 3.1; see Appendix B for the differences). We provide additional details about data preprocessing, sentence assignment, dataset splitting, and text formatting in Appendix A, with statistics of our datasets shown in Table 5. In the training portion, our CHILDES dataset contains 7,790 words to be meta-learned and has a total of $5.8$M tokens, while our BabyLM-10M dataset contains 15,821 words to be meta-learned and has a total of $7.8$M tokens. In comparison, a child receives roughly $3$M to $12$M words per year (Frank, 2023), and thus our training data is of a similar magnitude to a year’s worth of linguistic input for a child.
@@ 3.3 Held-out Word Classification
Given a query example sentence $q$ that uses a new word and a set of $C$ candidate words $\{w^{(c)}\}_{c=1}^{C}$, the task for the model is to match the query example to the most suitable one among the $C$ candidate words. Each $w^{(c)}$ is represented by a context containing a set of $K-1$ study examples $\{x^{(c)}_{k}\}_{k=1}^{K-1}$ illustrating its usage. (Note that in all query and study examples, the occurrences of the new word are replaced with the same special placeholder token, e.g., [new-token], as described in Section 3.1.) The context of $w^{(c)}$ is a sequence in the same format as the first $K-1$ examples in a training episode, ending with a separator token (e.g., <sep>): <sep> $x^{(c)}_{1}$ <sep> $\cdots$ <sep> $x^{(c)}_{K-1}$ <sep>. The query example is formatted as a continuation sequence of the context: $q$ <sep>. This formatting ensures that concatenating a context sequence and a query sequence results in a sequence with $K$ examples, just like a sequence for a meta-learning training episode. To determine the best match, we compute the conditional likelihood of the query sequence given the context: $p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The model predicts the word corresponding to the context with the highest likelihood: $\operatorname*{arg\,max}_{c}p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The prediction is correct if it is the ground-truth word in the query $q$.
We evaluate each model (trained from scratch or finetuned) by measuring the classification accuracy on held-out meta-learned words from the validation or test portions of the model’s training or finetuning corpus. For each evaluation, we group $C$ distinct meta-learned words into a $C$-way classification task. For each word, we sample $K-1$ study examples and one query example to construct the task. See Appendix C for an example task and additional details on task construction.
@@ 5.2 Syntactic Category Classification
Results
We first find that the Llama-3 8B baseline achieves 66% accuracy on this task, which is higher than random chance (50%), suggesting that it can infer the syntactic categories of new words in one shot and generalize them to novel contexts. The Minnow model improves Llama-3 8B’s accuracy to 83%, a 17% increase over the baseline. Meanwhile, both Minnow and CoLLEGe improve Llama-2 7B’s accuracy from 69% to 80%, which is an 11% increase. Fine-grained results from models finetuned with Minnow, trained from scratch, and CoLLEGe are provided in Appendix D. We find in all settings that the Minnow model improves accuracy by 4–23% compared to the baseline on all pairs of categories. These results show that Minnow finetuning effectively helps in learning the syntactic categories of new words and generalizing accordingly, and is comparable to CoLLEGe in improvements. In addition, note that our models are not specifically finetuned on this syntactic category classification task and dataset, demonstrating the generality of the acquired word learning ability.
5.3 New Usage Example Generation
New Usage Example Definition
Variant Method BabyLM-10M test Chimera CoLLEGe-DefGen
Llama-3 8B baseline 32 42 29
+Minnow 52 55 39
Llama-3 8B Instruct baseline 41 52 33
+Minnow 47 36 37
Llama-2 7B +CoLLEGe 13 33 4
+Minnow 74 58 47
Table 1: Percentages of wins of each model when comparing the generations from the pairs of models in each box, judged by GPT-4o. In the top two ruled rows, we compare Llama-3 8B baseline (pre-trained to instruction-tuned) with a Minnow model finetuned from that baseline. In the bottom-most ruled row, we compare CoLLEGe with the Minnow model finetuned from the pre-trained Llama-2 7B. Each new example or definition is generated by greedy decoding. (Results of top-p sampled generations are shown in Table 10 in Appendix E.) The percentage of ties is the remaining after subtracting the win percentages of the two models. GPT-4o more frequently chooses the Minnow model as the winner compared to the corresponding baseline model in all settings except for the instruction-tuned variant on Chimera.
@@ 5.3 New Usage Example Generation
The two tests we have described so far evaluate models in a discriminative setting. Here, we quantitatively and qualitatively evaluate if models use the new word appropriately in a generative setting. For a Minnow model finetuned with $K$ examples per episode, we evaluate it by showing it $K-1$ in-context study examples, formatted as a sequence in the classification setting (Section 3.3). We ask the model to do what it was trained for: We prompt the model with this sequence of study examples, which ends with a separator token, so the model will continue the sequence by generating a new usage example, ending with another separator token as End-Of-Sequence. For CoLLEGe, we generate a new example using the prompt “A single example sentence using the word ‘[new-token]’ (in one line):”.
We sample study examples from two datasets: the BabyLM-10M test portion in Section 3.2 and the Chimera dataset (Lazaridou et al., 2017). The Chimera dataset was specifically constructed for few-shot word learning. It has 33 different new words for learning, each referring to a “chimera” concept, i.e., a mixture of two existing and related concepts (e.g., cello and bagpipe). See Appendix F for more details of the dataset.
For the quantitative evaluation, we compare a pair of new usage examples generated from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. The comparison is simulated as a head-to-head competition following Teehan et al. (2024). Specifically, we provide GPT-4o (OpenAI, 2024) the same $K-1$ study examples in a list format with a pseudo-word “dax” as the placeholder for the word, as in the off-the-shelf baseline (without the last separator; Section 3.4.1), followed by a question “Which of the following is a better next example for the word ‘dax’, or they tie?” with three shuffled options, including the two generations and one “Tie”. (See Appendix E for prompting details.) The choice of GPT-4o decides whether and which one model wins the competition, or whether the models were tied in quality. For the qualitative evaluation, we manually pick meta-learned words (Table 2 and Appendix F) and examine the syntactic correctness and semantic appropriateness of the generated examples.
@@ Results
For the quantitative evaluation, Table 1 shows the percentages of wins of each of the baseline and the Minnow model on both the BabyLM-10M test portion and Chimera. Across all settings, the Minnow model wins more often than the corresponding baseline except for the instruction-tuned variant on Chimera, demonstrating the improvement brought by Minnow and its better performance compared to CoLLEGe. For the qualitative evaluation, Table 2 shows a word picked from the BabyLM-10M test portion along with its study and generated examples. See Appendix F for additional examples from the BabyLM-10M test portion and Chimera and detailed analysis of the generations. A manual analysis of these generated examples reveals that the Minnow model more often generates syntactically correct and semantically plausible new usage examples compared to the baseline, confirming that Minnow finetuning improves the ability to understand and use a new word. Nevertheless, in several cases, the Minnow model still shows obvious syntactic and factual errors and merely rewords the study examples.
@@ 5.4 Definition Generation
Table 3: Quantitative evaluation of 1-shot generated definitions by comparing them with ground-truth definitions. We sample an example per word from CoLLEGe-DefGen. All definitions are generated with greedy decoding. “FLAN-T5 XL +DefInstr” is a specialized definition-generation model from Giulianelli et al. (2023). “baseline” means using a pseudo-word ‘wug’ as the placeholder. Scores of Minnow models (“+Minnow”) are averaged across 3 runs. Finetuning Llama-3 8B with Minnow improves the baseline models on both datasets and both metrics, and the Minnow model finetuned from the instruction-tuned variant of Llama-3 8B performs the best. The instruction-tuned variant of Llama-3 8B (“Llama-3 8B Instruct”) is better than the pre-trained variant (“Llama-3 8B”) on definition generation likely due to its better instruction-following ability. The Minnow model also beats CoLLEGe except for the BERTScore F1 on CoLLEGe-DefGen. The specialized model (“FLAN-T5 XL +DefInstr”) performs the best on Oxford, but note that it is a model specialized in definition generation and was finetuned on Oxford.
@@
For the quantitative evaluation, we perform two types of comparison. The first type compares the model-generated and ground-truth definitions for each word by computing BERTScore F1 (Zhang et al., 2020) and ROUGE-L (Lin, 2004). The second type compares a pair of definitions generated from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. Similarly to what we did in Section 5.3, we ask GPT-4o a question (without usage examples): “Which of the following is a better definition for the word ‘Word’, or they tie?” where Word is the ground-truth word form, followed by three shuffled options including the two generated definitions and one “Tie” (see Appendix E for detailed prompting settings). [^9] We only perform this comparison on the CoLLEGe-DefGen dataset due to the large scale of the Oxford dataset. For the qualitative evaluation, we manually inspect 1-shot generated definitions for words from each dataset (presented in Table 4 and Tables 18 and 19 in Appendix G).
@@ Results
For the qualitative evaluation, Table 4 shows Minnow-model-generated and ground-truth definitions for a word from CoLLEGe-DefGen (see Tables 18 and 19 in Appendix G for additional examples from CoLLEGe-DefGen and Oxford). In our manual analysis, we find that definitions generated by the Minnow model often capture most of the word meanings, form reasonable inferences from the contexts, and outperform the baselines. However, they are not always precise compared to the ground-truth definitions.
@@ 6 Conclusion
The efficacy of Minnow, or meta-learning in general, suggests that human-level efficiency in linguistic generalizations may be acquired through practicing over many instances of learning tasks, without presuming strict, explicit inductive biases (Russin et al., 2024; Irie and Lake, 2024). Whether models achieve the generalizations in this work through human-like mechanisms remains for future analysis.
@@ 7 Limitations
In addition, we only consider learning a single new word on the fly, and each word is represented by the same special token. However, in real-world learning, both humans and models need to continually learn multiple words, usages, and even abstract rules (Sinha et al., 2023; Lampinen, 2024; Mueller et al., 2024). Future work could implement continual learning of multiple different words by meta-training to learn real words or pseudo-words. This way of meta-training could also allow the rapid word learning ability to co-exist with other general abilities of language models.
@@
In the paper, we focus on learning word-forms, each represented by a single special placeholder token, without considering morphological features. However, different words (word-forms) may be derived from the same root word and thus share a common underlying concept, but differ only in their affixes. Such a group of words forms a word family (Bauer and Nation, 1993), and such a family can decrease the time needed to derive and recognize related words (Nagy et al., 1989). In Appendix H, we discuss in detail our treatments of words and potential issues. In summary, our Minnow models are somehow robust to these variations. Future work could take morphological features into consideration during training, and conduct targeted evaluations of morphological variations of the learned words.
@@
In the generative evaluation settings (Section 5.3 and 5.4), we used ROUGE-L, BERTScore F1, and LLM-as-a-Judge (GPT-4o in our case) for automatic quantitative evaluations. However, these evaluations are imperfect. For example, different metrics are suitable for generations in different lengths, and LLM evaluators are known to have biases and inconsistencies (Doostmohammadi et al., 2024; Stureborg et al., 2024). Future work should conduct careful human evaluations for futher validation to avoid these potential issues.

### v3

@@ Abstract
Wentao Wang1    Guangyuan Jiang2    Tal Linzen1    Brenden M. Lake1 1New York University   2MIT {ww2135, linzen, brenden}@nyu.edujianggy@mit.edu
@@ 1 Introduction
We also finetune Llama-3 8B with Minnow to see if we can enhance its word-learning ability. In a series of discriminative and generative evaluations, we show that this improves Llama-3 8B’s ability to discriminate between new words, identify syntactic categories of new words, and generate reasonable new usages and definitions for new words, where each new word is learned from one or a few in-context examples. Most of these improvements are achieved without specific training on these evaluation tasks. We release our code at https://github.com/wwt17/meta-learning-word.
@@ 3.1 Method: Minnow
Note that our setting differs from previous MetaICL settings (Min et al., 2022; Chen et al., 2022; Lake and Baroni, 2023) in two ways. First, each example is not an input–output pair $(x^{(i)}_{k},y^{(i)}_{k})$, but just $x^{(i)}_{k}$. Second, there is no explicit separation between study examples and a query: [^3] The study examples (or support examples) are the small number of examples given for a new episode from which the model learns or adapts. A query example (or generalization example) is a new example on which the model is tested after learning from the study examples. These terms are used in few-shot meta-learning literature (e.g., Lake and Baroni, 2023). our setting effectively uses every example $x^{(i)}_{k}$ as a query with all previous examples $x^{(i)}_{1},\ldots,x^{(i)}_{k-1}$ as its study examples.
@@ 3.2 Data
Figure 2: An example task for held-out word classification. This is an example where we have $C=2$ candidate words and $K=4$ examples per word (left: three study examples per word; right: one query example per word). For each query example, we compute the conditional likelihood of its query sequence by prepending each context of study examples (lines in the middle), and we expect the correct context to give a higher likelihood (solid line) than the mismatched context (dashed line). See Appendix C for a full description.
@@
We split both the meta-learning component (by word) and the language modeling component (by sentence) into training (80%), validation (10%) and test (10%) portions. Each dataset is used for both training models from scratch and finetuning pre-trained Llama-3 8B, but the text is formatted and tokenized differently (in addition to the different special tokens in Section 3.1; see Appendix B for the differences). We provide additional details about data preprocessing, sentence assignment, dataset splitting, and text formatting in Appendix A, with statistics of our datasets shown in Table 6. In the training portion, our CHILDES dataset contains 7,790 words to be meta-learned and has a total of 5.85.8M tokens, while our BabyLM-10M dataset contains 15,821 words to be meta-learned and has a total of 7.87.8M tokens. In comparison, a child receives roughly 33M to 1212M words per year (Frank, 2023), and thus our training data is of a similar magnitude to a year’s worth of linguistic input for a child.
@@ 3.3 Held-out Word Classification
Given a query example sentence $q$ that uses a new word and a set of $C$ candidate words $\{w^{(c)}\}_{c=1}^{C}$, the task is to use the model likelihoods to match the query example to the most suitable one among the $C$ candidate words. Each $w^{(c)}$ is represented by a context containing a set of K1K-1 study examples $\{x^{(c)}_{k}\}_{k=1}^{K-1}$ illustrating its usage. (Note that in all query and study examples, the occurrences of the new word are replaced with the same special placeholder token, e.g., [new-token], as described in Section 3.1.) The context of $w^{(c)}$ is a sequence in the same format as the first K1K-1 examples in a training episode, ending with a separator token (e.g., <sep>): <sep> $x^{(c)}_{1}$ <sep> $\cdots$ <sep> $x^{(c)}_{K-1}$ <sep>. The query example is formatted as a continuation sequence of the context: $q$ <sep>. This formatting ensures that concatenating a context sequence and a query sequence results in a sequence with $K$ examples, just like a sequence for a meta-learning training episode. To determine the best match, we compute the conditional likelihood of the query sequence given the context: $p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. We then choose the word among the $C$ candidate words that gives the highest likelihood: $\operatorname*{arg\,max}_{c}p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The choice is correct if it is the ground-truth word in the query $q$.
We evaluate each model (trained from scratch or finetuned) by measuring the classification accuracy on held-out meta-learned words from the validation or test portions of the model’s training or finetuning corpus. For each evaluation, we group $C$ distinct meta-learned words into a $C$-way classification task. For each word, we sample K1K-1 study examples and one query example to construct the task. Figure 2 shows an example task. See Appendix C for additional details on task construction.
@@ 5.2 Syntactic Category Classification
Variant Method Mean Acc. (%)
from scratch Minnow 77
Llama-3 8B baseline 66
+Minnow 83
Llama-2 7B baseline 69
+CoLLEGe 80
+Minnow 80
Table 1: Mean accuracies of each model on the syntactic category classification task. The random chance level accuracy is 50%. See Appendix D for fine-grained results. In the top row, we show the result of the model trained from scratch with Minnow on BabyLM-10M (Section 4). In the other two ruled rows, we show models based on Llama-3 8B and Llama-2 7B, respectively. We only have results for CoLLEGe based on Llama-2 7B because the original checkpoint is based on Llama-2 7B. Minnow accuracies are much higher than random chance and their corresponding baselines.
@@ Results
Mean accuracies on this syntactic category classification task are summarized in Table 1. We first find that the Llama-3 8B baseline achieves 66% accuracy on this task, which is higher than random chance (50%), suggesting that it can infer the syntactic categories of new words in one shot and generalize them to novel contexts. The Minnow model improves Llama-3 8B’s accuracy to 83%, an increase of 17% points over the baseline. Meanwhile, both Minnow and CoLLEGe improve Llama-2 7B’s accuracy from 69% to 80%, which is an increase of 11% points. Fine-grained results from models finetuned with Minnow, trained from scratch, and CoLLEGe are provided in Appendix D. We find in all settings that the Minnow model improves accuracy by 4–23% compared to the baseline on all pairs of categories. These results show that Minnow finetuning effectively helps in learning the syntactic categories of new words and generalizing accordingly, and is comparable to CoLLEGe in improvements. In addition, note that our models are not specifically finetuned on this syntactic category classification task and dataset, demonstrating the generality of the acquired word learning ability.
5.3 New Usage Example Generation
The two tests we have described so far evaluate models in a discriminative setting. Here, we quantitatively and qualitatively evaluate if models use the new word appropriately in a generative setting. For a Minnow model finetuned with $K$ examples per episode, we evaluate it by showing it K1K-1 in-context study examples, formatted as a sequence in the classification setting (Section 3.3). We ask the model to do what it was trained for: We prompt the model with this sequence of study examples, which ends with a separator token, so the model will continue the sequence by generating a new usage example, ending with another separator token as End-Of-Sequence. For CoLLEGe, we generate a new example using the prompt “A single example sentence using the word ‘[new-token]’ (in one line):”.
We sample study examples from two datasets: the BabyLM-10M test portion in Section 3.2 and the Chimera dataset (Lazaridou et al., 2017). The Chimera dataset was specifically constructed for few-shot word learning. It has 33 different new words for learning, each referring to a “chimera” concept, i.e., a mixture of two existing and related concepts (e.g., cello and bagpipe). The usage examples of a new word are sentences using one of the components of the chimera, randomly extracted from a large corpus. See Appendix F for more details of the dataset.
For the quantitative evaluation, we compare a pair of new usage examples generated from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. The comparison is simulated as a head-to-head competition following Teehan et al. (2024). Specifically, we provide GPT-4o (OpenAI, 2024) the same K1K-1 study examples in a list format with a pseudo-word “dax” as the placeholder for the word, as in the off-the-shelf baseline (without the last separator; Section 3.4.1), followed by a question “Which of the following is a better next example for the word ‘dax’, or they tie?” with three shuffled options, including the two generations and one “Tie”. (See Appendix E for prompting details.) The choice of GPT-4o decides whether and which one model wins the competition, or whether the models were tied in quality. For the qualitative evaluation, we manually pick meta-learned words (Table 2 and Appendix F) and examine the syntactic correctness and semantic appropriateness of the generated examples.
New Usage Example Definition
Variant Method BabyLM-10M test Chimera CoLLEGe-DefGen
Llama-3 8B baseline 31 53 27
+Minnow 50 42 40
Llama-3 8B Instruct baseline 40 44 32
+Minnow 46 48 37
Llama-2 7B +CoLLEGe 14 29 4
+Minnow 73 63 46
Table 3: Percentages of wins of each model when comparing the generations from the pairs of models in each box, judged by GPT-4o. In the top two ruled rows, we compare Llama-3 8B baseline (pre-trained to instruction-tuned) with a Minnow model finetuned from that baseline (averaged across 3 runs). In the bottom-most ruled row, we compare CoLLEGe with the Minnow model finetuned from the pre-trained Llama-2 7B (averaged across 3 runs). The left two datasets are for new usage example generation (Section 5.3; each new usage example is generated by providing 4 study examples), and the right-most one is for definition generation (Section 5.4; each definition is generated by providing 3 study examples). Each new example or definition is generated by greedy decoding. We boldface significantly more preferred models ($p<.05$ in paired t-tests across 3 runs). (Results of top-p sampled generations are shown in Table 12 in Appendix E.) The percentage of ties is the remaining after subtracting the win percentages of the two models. On average, GPT-4o more frequently chooses the Minnow model as the winner compared to the corresponding baseline model in all settings except for the pretrained Llama-3 8B on Chimera. The improvements by Minnow on the instruction-tuned Llama-3 8B are not significant ($p>.1$).
Results
@@
For the quantitative evaluation, Table 3 shows the percentages of wins of each of the baseline and the Minnow model on both the BabyLM-10M test portion and Chimera. Across all settings, the Minnow model wins more often on average than the corresponding baseline except for the pretrained Llama-3 8B on Chimera, demonstrating the improvement brought by Minnow and its better performance compared to CoLLEGe. For the qualitative evaluation, Table 2 shows a word picked from the BabyLM-10M test portion along with its study and generated examples. See Appendix F for additional examples from the BabyLM-10M test portion and Chimera and detailed analysis of the generations. A manual analysis of these generated examples reveals that the Minnow model more often generates syntactically correct and semantically plausible new usage examples compared to the baseline, confirming that Minnow finetuning improves the ability to understand and use a new word. Nevertheless, in several cases, the Minnow model still shows obvious syntactic and factual errors and merely rewords the study examples.
@@ 5.4 Definition Generation
Table 4: Quantitative evaluation of 1-shot generated definitions by comparing them with ground-truth definitions. See Table 23 in Appendix G for results from all models. We sample an example per word from CoLLEGe-DefGen. All definitions are generated with greedy decoding. “FLAN-T5 XL +DefInstr” is a specialized definition-generation model from Giulianelli et al. (2023). “baseline” means using a pseudo-word ‘wug’ as the placeholder. Scores of Minnow models (“+Minnow”) are averaged across three runs. Finetuning Llama-3 8B with Minnow improves the baseline models on both datasets and both metrics ($p<.01$), and the Minnow model finetuned from the instruction-tuned variant of Llama-3 8B performs the best on CoLLEGe-DefGen ($p<.01$; likely due to its better instruction-following ability; no significant difference is found on Oxford). The Minnow model beats CoLLEGe on ROUGE-L ($p=.027$ on CoLLEGe-DefGen and $p=.012$ on Oxford) but not on BERTScore F1 ($p=.289$ on Oxford). The specialized model (“FLAN-T5 XL +DefInstr”) performs the best on Oxford ($p<.01$), but note that it is specialized in definition generation and was finetuned on Oxford.
@@
For the quantitative evaluation, we perform two types of comparison. The first type compares the model-generated and ground-truth definitions for each word by computing BERTScore F1 (Zhang et al., 2020) and ROUGE-L (Lin, 2004). The second type compares a pair of definitions generated from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. Similarly to what we did in Section 5.3, we ask GPT-4o a question (without usage examples): “Which of the following is a better definition for the word ‘Word’, or they tie?” where Word is the ground-truth word form, followed by three shuffled options including the two generated definitions and one “Tie” (see Appendix E for detailed prompting settings). [^11] We only perform this comparison on the CoLLEGe-DefGen dataset due to the large scale of the Oxford dataset. For the qualitative evaluation, we manually inspect 1-shot generated definitions for words from each dataset (presented in Table 5 and Tables 25 and 26 in Appendix G).
@@ Results
For the qualitative evaluation, Table 5 shows Minnow-model-generated and ground-truth definitions for a word from CoLLEGe-DefGen (see Tables 25 and 26 in Appendix G for additional examples from CoLLEGe-DefGen and Oxford). In our manual analysis, we find that definitions generated by the Minnow model often capture most of the word meanings, form reasonable inferences from the contexts, and outperform the baselines. However, they are not always precise compared to the ground-truth definitions.
@@ 6 Conclusion
The efficacy of Minnow, or meta-learning in general, suggests that human-level efficiency in linguistic generalizations may be acquired through practicing over many instances of learning tasks, without presuming strict, explicit inductive biases (Russin et al., 2024; Irie and Lake, 2024). Whether models achieve the generalizations in this work through human-like mechanisms, such as systematicity and categorical abstraction, remains for future analysis.
@@ 7 Limitations
In addition, we only consider learning a single new word on the fly, and each word is represented by the same special token. However, in real-world learning, both humans and models need to continually learn multiple words, usages, and even abstract rules (Sinha et al., 2023; Lampinen, 2024; Mueller et al., 2024). Future work could implement continual learning of multiple different words by meta-training to learn real words or pseudo-words, and by utilizing long-term memory. In that way, we could also allow the rapid word learning ability to co-exist with other general abilities of language models.
@@
In the paper, we focus on learning word-forms, each represented by a single special placeholder token, without considering morphological features. However, morphological features in a word may allow systematic recombination of meanings from related words, decrease the time needed to derive and recognize the word (Nagy et al., 1989), and help children to learn the word (Moore and Bergelson, 2024). In Appendix H, we discuss in detail our treatments of words and potential issues. In summary, our Minnow models are somehow robust to these variations. Future work could take morphological features into consideration during training, and conduct targeted evaluations of morphological variations of the learned words.
@@
In the generative evaluation settings (Section 5.3 and 5.4), we used ROUGE-L, BERTScore F1, and LLM-as-a-Judge (GPT-4o in our case) for automatic quantitative evaluations. However, these evaluations are imperfect. For example, different metrics are suitable for generations in different lengths, and LLM evaluators are known to have biases and inconsistencies (Doostmohammadi et al., 2024; Stureborg et al., 2024). Future work should conduct careful human evaluations for further validation to avoid these potential issues.